A food company has developed a high mineral sea salt (sodium). A nurse practitioner hypothesizes that getting more sodium from the new sea salt predicts an increase in blood pressure. Below are the sodium and BP measurements for a sample of participants that regularly use the new sea salt. What can the nurse practitioner conclude with an α of 0.10?

Sodium BP
8.3
8.2
8.3
8.2
8.4
8.4
8.3
8.2
8.3
167
196
190
187
149
141
160
190
175

a) What is the appropriate statistic?
---Select one--- (na, Correlation, Slope, Chi-Square)
Compute the option selected in a):  

b) Obtain/compute the appropriate values to make a decision about H0.
(Hint: Make sure to write down the null and alternative hypotheses to help solve the problem.)
Critical value =  ; Test statistic =  
Decision:  ---Select one--- (Reject H0, Fail to reject H0)

d) Make an interpretation based on the results.

a. More sodium intake significantly predicts an increase in blood pressure.

b. More sodium intake significantly predicts a decrease in blood pressure.

c. Sodium intake does not significantly predict blood pressure.
Answer #1

(a) The appropriate statistic is Correlation

To calculate pearsons co efficient r


From the data given, we get the following table and values

Count x y xy x2 y2
1 8.3 167 1386.1 68.89 27889
2 8.2 196 1607.2 67.24 38416
3 8.3 190 1577 68.89 36100
4 8.2 187 1533.4 67.24 34969
5 8.4 149 1251.6 70.56 22201
6 8.4 141 1184.4 70.56 19881
7 8.3 160 1328 68.89 25600
8 8.2 190 1558 67.24 36100
9 8.3 175 1452.5 68.89 30625
Total 74.6 1555 12878.2 618.4 271781

SUM (x) = 74.6, SUM (y) = 1555, SUM (xy) = 12878.2,

SUM (x2) = 618.4, (SUM x)2 = (74.6)2= 5565.16

SUM (y2) = 271781, (SUM y)2 = (1555)2 = 2418025

SUM (x) x SUM (y) = 74.6 x 1555 = 116003

n = 9

r = -0.8937

_________________________________________

(b) The Hypothesis Test

The Hypothesis:

H0: = 0 : There is no linear relationship between the 2 variables.

Ha : 0 : There is a linear relationship between the 2 variables.

The Test Statistic:

The Critical Value (2 tail): at = 0.05, df = n -2 = 7 is -2.365 and +2.365

The Decision Rule: If T obtained is < -t critical or if t obtained is > t critical, then Reject H0.

The Decision: Since t obtained (-5.27) is < -t critical(-2.365), We Reject H0.

The Conclusion: There is sufficient evidence at the = 0.01 level to conclude that there is a linear relationship between the 2 variables (i.e r is significant).

_________________________________________________________

Interpretation. Since there is a significant relationship and the same is negative, therefore

Option b: More sodium intake significantly predicts a decrease in blood pressure.

__________________________________________________________